# Uncertainty Quantification for Retrieval-Augmented Reasoning

### v1

@@ 5.1. Direct Evaluation of UQ Estimations
Evaluation Metrics: To evaluate the quality of outputs, we follow Jin et al. (2025) and report the exact match, where a prediction is counted as correct if and only if it exactly matches one of the ground-truth responses. For evaluating UQ methods, we follow prior work on UQ and use the threshold-free metric AUROC, which captures the correlation between uncertainty scores and response correctness (Kadavath et al., 2022; Kuhn et al., 2023; Bakman et al., 2024). As suggested by Perez-Beltrachini and Lapata (2025), significant differences between two AUROC values are assessed using the paired De Long test (DeLong et al., 1988).
Models: In line with prior work in RAR (Jin et al., 2025; Chen et al., 2025c), we employ Qwen-2.5-7B-Instruct (Yang et al., 2024) as the generator LLM and path summary generator for action $a_{av}$. For UQ, we sample $10$ responses per query with a temperature of $T=1.0$, while for correctness evaluation we generate the most-likely generation with $T=0.7$ (Soudani et al., 2025; Bakman et al., 2024). Retrieval is performed using a two-stage re-ranking pipeline: BM25 (Robertson and Walker, 1994) is used for initial retrieval, followed by re-ranking with the pre-trained cross-encoder model ms-marco-MiniLM-L-6-v2 from the sentence-transformers library. [^4] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 All experiments are conducted on four Nvidia A100 GPU, each with 40 GB memory, requiring $\sim$1500 GPU hours in total.
Baselines: We use two sets of baselines: (1) Path-based methods, which focus on generating multiple responses based on diverse reasoning paths and differ mainly in how they initiate new generations relative to the most-likely generation. Self-Consistency (SelfC) (Wang et al., 2023) ignores the most-likely generation and instead produces diverse set of reasoning paths, starting from scratch. Reasoning Consistency (ReaC) (Qi et al., 2025) randomly truncates the most-likely reasoning path at different random steps and regenerates the response based on the subsequent reasoning steps. Retrieval-Retained Reasoning Consistency (RrrC) (Li et al., 2025b) applies truncation only after the last retrieved document. In all cases, the generated responses are aggregated into a final score using majority voting. 2) Estimation-based methods include both white-box and black-box approaches. The white-box methods are PE (Kadavath et al., 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), LARS (Yaldiz et al., 2025a), and SAR (Duan et al., 2024). The black-box methods are NumSS (Kuhn et al., 2023), EigV, ECC, Deg (Lin et al., 2024), and P(true) (Kadavath et al., 2022). All of these methods rely on generations obtained in the same way as SelfC. All UQ methods are implemented using TruthTorchLM (Yaldiz et al., 2025b).
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025a) ReSearch (Chen et al., 2025c) Search-R1 (Jin et al., 2025) Avg.
@@
RrrC (Li et al., 2025b) 61.4 65.6 73.0 60.8 72.2 82.4 56.6 66.2 68.2 59.2 64.4 64.0 60.6 64.0 63.6 65.48
P(true) (Kadavath et al., 2022) 70.6 70.8 80.4 70.4 67.2 75.0 72.6 65.8 75.0 78.8 71.4 78.2 70.8 70.0 74.4 72.76
@@
RrrC (Li et al., 2025b) 62.52 70.54 82.75 71.59 80.82 89.69 56.33 71.50 78.99 54.86 63.37 73.68 55.12 61.20 73.92 69.79
P(true) (Kadavath et al., 2022) 73.60 75.33 88.27 72.07 73.46 84.58 76.50 71.73 84.51 81.20 73.46 85.82 71.03 71.15 83.37 77.74
@@ 5.2. Extrinsic UQ Evaluation via Abstention
Task Formulation. Declining to respond due to uncertainty is an important application of UQ (Feng et al., 2024). An abstention function determines whether the model should withhold an answer. In our setup, this decision is guided by uncertainty scores. We introduce a threshold $\tau_{\text{abs}}$: if the uncertainty exceeds $\tau_{\text{abs}}$, the model abstains; otherwise, it produces an answer. Formally, given a response $r$ to a query $x$ and an uncertainty function $U$, we define the abstention function $f_{\text{abs}}$ as:
$f_{\text{abs}}(r)=\begin{cases}\text{true},&\text{if }U(x,r)>\tau_{\text{abs}}\\ \text{false},&\text{otherwise}.\end{cases}$
Baselines & Evaluation Metrics. As baselines, we implement the abstention task using uncertainty scores derived from different UQ methods, including P(true), RrrC, ReaC, SelfC, and $\text{R}^{2}$C. For evaluation, we follow Feng et al. (2024) and report results on AbstainAccuracy and AbstainF1. AbstainAccuracy evaluates whether the abstain decisions are correct, and AbstainF1 captures the balance between reliability and answer coverage. The evaluation metrics are explained in detail in Appendix B.1. The threshold $\tau_{\text{abs}}$ for each UQ method is selected by performing sweep using validation sets; see Appendix B.2 for more details.
@@ 5.3. Extrinsic UQ Evaluation via Model Selection
Task Formulation. Model selection (or selection-based model ensemble (Chen et al., 2025b; Jiang et al., 2023a)) aims to select a final response for a question based on multiple candidate responses generated by different systems (Chen et al., 2025b). Formally, given a user query $x$ and a set of systems $\{S_{1},S_{2},\dots,S_{M}\}$, each system $S_{i}$ produces a response candidate $r_{i}$. A model selection method $\mathcal{M}(x,R)$ then selects the final response $\hat{r}$ from all candidate responses $R=\{r_{1},r_{2},\dots,r_{M}\}$.
$\text{R}^{2}$C Select. The proposed $\text{R}^{2}$C Select utilizes the uncertainty scores derived from $\text{R}^{2}$C. Our method groups semantically similar responses into $K$ clusters. Following (Li et al., 2025b; Yaldiz et al., 2025b), we use Qwen-2.5-7B-Instruct to compute pairwise semantic similarities between candidate responses and cluster similar ones together. We then assign an uncertainty score to each cluster $c_{i}$:
$R_{C}=\{\langle r_{{}_{1}},u_{c_{1}}\rangle,\langle r_{{}_{2}},u_{c_{1}}\rangle,\dots,\langle r_{{}_{K}},u_{c_{K}}\rangle\}.$
The uncertainty score of each cluster $c=\{r_{i}\}_{i=1}^{m}$ is computed by aggregating the uncertainty scores of its members: $u_{c}=\sum_{r_{i}\in c}U(x,r_{i}),$ where the $U$ function provides the uncertainty score for the response $r_{i}$. The final response $\hat{r}$ is the response with the lowest uncertainty score: $\hat{r}=\arg\min_{j}u_{c_{j}}.$ If no clustering is applied, each response constitutes a cluster and $M=K$; we refer to this variation as $\text{R}^{2}$C Select w/o clustering.
Baselines & Evaluation Metrics. We evaluate $\text{R}^{2}$C Select performance against both single RAG and RAR systems, as well as existing selection-based ensemble approaches. As baselines, we consider LLMBlender (Jiang et al., 2023a) and RAGEnsemble (Chen et al., 2025a). LLMBlender is a trained reward model that given a user query and a set of candidate responses, ranks the responses accordingly. RAGEnsemble is an instruction-based approach that selects a single final answer from a set of candidate responses. For all selection methods, including the baselines and $\text{R}^{2}$C Select, the response candidates are obtained from SelfAsk (Press et al., 2023), ReAct (Yao et al., 2023), Search-o1 (Li et al., 2025a), ReSearch (Chen et al., 2025c), and Search-R1 (Jin et al., 2025). For evaluation, we report the correctness of the final answer using exact match.
6. Results
We present a set experiments that address the following research questions: RQ1: How does $\text{R}^{2}$C perform in quantifying uncertainty for different RAR models? (Sec. 6.1), RQ2: How does $\text{R}^{2}$C perform as an external signal on downstream tasks, such as Abstention and Model Selection? (Sec. 6.2), RQ3: What factors contribute to the effectiveness of $\text{R}^{2}$C? (Sec. 6.3), RQ4: How does $\text{R}^{2}$C balance effectiveness and efficiency? (Sec. 6.4), RQ5: What is the effect of different actions in the performance of $\text{R}^{2}$C? (Sec. 6.5).
6.1. Uncertainty Quantification Performance
RQ1 evaluates the performance of $\text{R}^{2}$C compared to other UQ methods. Table 1 presents results on five RAR systems across three datasets. The findings indicate that most white-box methods, i.e., PE (Kadavath et al., 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), and SAR (Duan et al., 2024), perform relatively poorly, with AUROC values ranging from about 30 to 60. This weakness stems from their overreliance on token probabilities. In contrast, black-box methods, such as NumSS (Kuhn et al., 2023), EigV, ECC, and Deg (Lin et al., 2024), generally outperform white-box methods, reaching average AUROC values between roughly 60 and 80. Interestingly, P(true) ranks as the second-best method in terms of average AUROC across all approaches, highlighting that black-box methods become much more effective when using an LLM as the scoring function. This advantage largely comes from their stronger reliance on textual diversity. The supervised method LARS, optimized for QA with no reasoning, achieves AUROC scores between 65 and 85 across all cases, surpassing both white-box and black-box approaches on average. This finding highlights the potential of as a promising direction for future research.
@@ 6.1. Uncertainty Quantification Performance
Search-o1 (Li et al., 2025a) 33.2 29.0 10.0 24.1
ReSearch (Chen et al., 2025c) 38.6 38.8 16.6 31.3
@@
- w/o clus. & unc. 32.0 26.0 8.6 22.2
RAGEnsemble (Chen et al., 2025a) 45.6 46.0 19.4 37.0
@@ 6.2. Extrinsic Evaluation Results
For model selection, Table 3 reports the results of individual RAG and RAR systems, as well as selection-based ensemble models. Among individual systems, RAR models outperform RAG models for both simpler single-hop questions (PopQA) and more complex multi-hop queries (HotpotQA and Musique). Within selection-based ensemble models, LLMBLender performs poorly, even worse than fine-tuned single RAR models such as ReSearch and Search-R1. In contrast, RAGEnsemble surpasses both LLMBLender and all single RAR models, highlighting the advantage of an instruction-tuned model for selecting the final response. Finally, $\text{R}^{2}$C Select achieves the best performance, significantly outperforming all RAR and RAG systems, with 3.7% average improvement on HotpotQA and Musique datasets. These results confirm that $\text{R}^{2}$C score is a reliable criterion, not only within a single system but also across different systems. Interestingly, using $\text{R}^{2}$C uncertainty scores alone, without clustering, improves selection-based ensemble models in comparable settings, highlighting informativeness of $\text{R}^{2}$C scores even for models that are not trained on uncertainty.
@@ 6.3. Explaining the Strength of $\text{R}^{2}$C
RQ3 examines why $\text{R}^{2}$C outperforms other UQ methods. To this end, we analyze the diversity of reasoning paths as an indicator of how effectively $\text{R}^{2}$C captures the uncertainty arising from both the generator and the retriever, i.e., the two key sources of uncertainty in our framework. We report the number of unique documents retrieved during the multi-generation step, as shown in Figure 2. On average, $\text{R}^{2}$C retrieves 24.71 unique documents for a single uncertainty score, whereas RrrC, SelfC, and ReaC retrieve 5.81, 15.35, and 16.4 documents, respectively.
@@
where $cos(.)$ represents cosine similarity function, and $e_{i}$ denotes the normalized embedding of query $q_{i}$ (i.e., $|e_{i}|=1$) obtained from sentence-transformers/all-MiniLM-L6-v2. Figure 3 presents the results. On average, $\text{R}^{2}$C achieves a diversity score of $0.35$, while RrrC, SelfC, and ReaC achieve $0.20$, $0.28$, and $0.30$, respectively. These results indicate that $\text{R}^{2}$C generates more diverse search queries when estimating uncertainty. These findings indicate $\text{R}^{2}$C effectively captures the uncertainty of both the retriever and the generator by sequentially diversifying their inputs.
@@
For a deeper exploration of the trade-offs between effectiveness and efficiency, we measure efficiency using the total number of generated tokens, following prior work (Perez-Beltrachini and Lapata, 2025). Figure 5 reports AUROC performance for equal numbers of generated tokens. It shows $\text{R}^{2}$C achieves higher AUROC scores than other methods given the same number of generated tokens for multi-hop datasets. For the single-hop dataset, PopQA, $\text{R}^{2}$C performs comparably to SelfAsk, ReAct, and ReSearch, but surpasses them on Search-o1 and Search-R1. Moreover, $\text{R}^{2}$C produces on average around 700 tokens with three generations, reaching the same AUROC score of baselines with approximately 1,700 tokens with 10 generations. These findings demonstrate that $\text{R}^{2}$C improves efficiency by roughly 2.5 times. Overall, these results indicate that $\text{R}^{2}$C outperforms other UQ methods in terms of both effectiveness and token generation efficiency.
Figure 6. Performance of $\text{R}^{2}$C with different action sets.

### v2

@@ 5.1. Direct Evaluation of UQ Estimations
Evaluation Metrics: To evaluate the quality of outputs, we follow Jin et al. (2025) and report the exact match, where a prediction is counted as correct if and only if it exactly matches one of the ground-truth responses. For evaluating UQ methods, we follow prior work on UQ and use the threshold-free metric AUROC, which captures the correlation between uncertainty scores and response correctness (Kadavath and others, 2022; Kuhn et al., 2023; Bakman et al., 2024). As suggested by Perez-Beltrachini and Lapata (2025), significant differences between two AUROC values are assessed using the paired De Long test (DeLong et al., 1988).
@@
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) ReSearch (Chen et al., 2025a) Search-R1 (Jin et al., 2025) Avg.
@@
RrrC (Li et al., 2025a) 61.4 65.6 73.0 60.8 72.2 82.4 56.6 66.2 68.2 59.2 64.4 64.0 60.6 64.0 63.6 65.48
P(true) (Kadavath and others, 2022) 70.6 70.8 80.4 70.4 67.2 75.0 72.6 65.8 75.0 78.8 71.4 78.2 70.8 70.0 74.4 72.76
@@
RrrC (Li et al., 2025a) 62.52 70.54 82.75 71.59 80.82 89.69 56.33 71.50 78.99 54.86 63.37 73.68 55.12 61.20 73.92 69.79
P(true) (Kadavath and others, 2022) 73.60 75.33 88.27 72.07 73.46 84.58 76.50 71.73 84.51 81.20 73.46 85.82 71.03 71.15 83.37 77.74
@@
Models: In line with prior work in RAR (Jin et al., 2025; Chen et al., 2025a), we employ Qwen-2.5-7B-Instruct (Yang and others, 2024) as the generator LLM and path summary generator for action $a_{av}$. For UQ, we sample $10$ responses per query with a temperature of $T=1.0$, while for correctness evaluation we generate the most-likely generation with $T=0.7$ (Soudani et al., 2025; Bakman et al., 2024). Retrieval is performed using a two-stage re-ranking pipeline: BM25 (Robertson and Walker, 1994) is used for initial retrieval, followed by re-ranking with the pre-trained cross-encoder model ms-marco-MiniLM-L-6-v2 from the sentence-transformers library. [^5] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L6-v2 All experiments are conducted on four Nvidia A100 GPU, each with 40 GB memory, requiring $\sim$1500 GPU hours in total.
Baselines: We use two sets of baselines: (1) Path-based methods, which focus on generating multiple responses based on diverse reasoning paths and differ mainly in how they initiate new generations relative to the most-likely generation. Self-Consistency (SelfC) (Wang et al., 2023) ignores the most-likely generation and instead produces diverse set of reasoning paths, starting from scratch. Reasoning Consistency (ReaC) (Qi et al., 2025) randomly truncates the most-likely reasoning path at different random steps and regenerates the response based on the subsequent reasoning steps. Retrieval-Retained Reasoning Consistency (RrrC) (Li et al., 2025a) applies truncation only after the last retrieved document. In all cases, the generated responses are aggregated into a final score using majority voting. 2) Estimation-based methods include both white-box and black-box approaches. The white-box methods are PE (Kadavath and others, 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), LARS (Yaldiz et al., 2025a), and SAR (Duan et al., 2024). The black-box methods are NumSS (Kuhn et al., 2023), EigV, ECC, Deg (Lin et al., 2024), and P(true) (Kadavath and others, 2022). All of these methods rely on generations obtained in the same way as SelfC. All UQ methods are implemented using TruthTorchLM (Yaldiz et al., 2025b).
@@ 5.2. Extrinsic UQ Evaluation via Abstention
Task Formulation. Declining to respond due to uncertainty is an important application of UQ (Feng et al., 2024). An abstention function determines whether the model should withhold an answer. In our setup, this decision is guided by confidence scores. We introduce a threshold $\tau_{\text{abs}}$: if the confidence is less than $\tau_{\text{abs}}$, the model abstains; otherwise, it produces an answer. Formally, given a response $r$ to a query $x$ and a confidence function $C$, we define the abstention function $f_{\text{abs}}$ as:
$f_{\text{abs}}(r)=\begin{cases}\text{true},&\text{if }C(x,r)<\tau_{\text{abs}}\\ \text{false},&\text{otherwise}.\end{cases}$
Baselines & Evaluation Metrics. As baselines, we implement the abstention task using confidence scores derived from different UQ methods, including P(true), RrrC, ReaC, SelfC, and $\text{R}^{2}$C. For evaluation, we follow Feng et al. (2024) and report two metrics: AbstainAccuracy and AbstainF1. These metrics assess how well a model balances answer correctness with appropriate abstention. We define a confusion matrix with four outcomes: (A) answered correctly, (B) abstained correctly, (C) answered incorrectly, and (D) abstained incorrectly. Based on this matrix, AbstainAccuracy is defined as $\frac{A+D}{A+B+C+D}$ and measures whether abstention decisions are correct overall. Ideally, a model should abstain when it would otherwise answer incorrectly, and provide an answer when it is correct. AbstainF1 captures the trade-off between reliability and answer coverage. It is computed as the harmonic mean of precision ($\frac{D}{B+D}$) and recall ($\frac{D}{C+D}$). This metric penalizes both unnecessary abstentions and incorrect answers, providing a balanced evaluation of abstention behavior. The threshold $\tau_{\text{abs}}$ for each UQ method is selected by performing sweep using validation sets.
@@ 5.3. Extrinsic UQ Evaluation via Model Selection
Task Formulation. Model selection (or selection-based model ensemble (Chen et al., 2025c; Jiang et al., 2023a)) aims to select a final response for a question based on multiple candidate responses generated by different systems (Chen et al., 2025c). Formally, given a user query $x$ and a set of systems $\{S_{1},S_{2},\dots,S_{M}\}$, each system $S_{i}$ produces a response candidate $r_{i}$. A model selection method $\mathcal{M}(x,R)$ then selects the final response $\hat{r}$ from all candidate responses $R=\{r_{1},r_{2},\dots,r_{M}\}$.
$\text{R}^{2}$C Select. The proposed $\text{R}^{2}$C Select utilizes the confidence scores derived from $\text{R}^{2}$C. Our method groups semantically similar responses into $K$ clusters. Following (Li et al., 2025a; Yaldiz et al., 2025b), we use Qwen-2.5-7B-Instruct to compute pairwise semantic similarities between candidate responses and cluster similar ones together. We then assign a confidence score to each cluster $g_{i}$:
$R_{G}=\{\langle r_{{}_{1}},c_{g_{1}}\rangle,\langle r_{{}_{2}},c_{g_{1}}\rangle,\dots,\langle r_{{}_{K}},c_{g_{K}}\rangle\}.$
The confidence score of each cluster $g=\{r_{i}\}_{i=1}^{m}$ is computed by aggregating the confidence scores of its members: $c_{g}=\sum_{r_{i}\in g}C(x,r_{i}),$ where the $C$ function provides the confidence score for the response $r_{i}$. The final response $\hat{r}$ is the response with the highest confidence score: $\hat{r}=\arg\max_{j}c_{g_{j}}.$ If no clustering is applied, each response constitutes a cluster and $M=K$; we refer to this variation as $\text{R}^{2}$C Select w/o clustering.
@@
Search-o1 (Li et al., 2025b) 33.2 29.0 10.0 24.1
ReSearch (Chen et al., 2025a) 38.6 38.8 16.6 31.3
@@
- w/o clus. & conf. 32.0 26.0 8.6 22.2
RAGEnsemble (Chen et al., 2025b) 45.6 46.0 19.4 37.0
@@
Baselines & Evaluation Metrics. We evaluate $\text{R}^{2}$C Select performance against both single RAG and RAR systems, as well as existing selection-based ensemble approaches. As baselines, we consider LLMBlender (Jiang et al., 2023a) and RAGEnsemble (Chen et al., 2025b). LLMBlender is a trained reward model that given a user query and a set of candidate responses, ranks the responses accordingly. RAGEnsemble is an instruction-based approach that selects a single final answer from a set of candidate responses. For all selection methods, including the baselines and $\text{R}^{2}$C Select, the response candidates are obtained from SelfAsk (Press et al., 2023), ReAct (Yao et al., 2023), Search-o1 (Li et al., 2025b), ReSearch (Chen et al., 2025a), and Search-R1 (Jin et al., 2025). For evaluation, we report the correctness of the final answer using exact match.
6. Results
We present a set experiments that address the following research questions: RQ1: How does $\text{R}^{2}$C perform in quantifying uncertainty for different RAR models? (Sec. 6.1), RQ2: How does $\text{R}^{2}$C perform as an external signal on downstream tasks, such as Abstention and Model Selection? (Sec. 6.2), RQ3: What factors contribute to the effectiveness of $\text{R}^{2}$C? (Sec. 6.3), RQ4: How does $\text{R}^{2}$C balance effectiveness and efficiency? (Sec. 6.4), RQ5: What is the effect of different actions in the performance of $\text{R}^{2}$C? (Sec. 6.5). RQ6: Does $\text{R}^{2}$C generalize across different backbone LLMs? (Sec. 6.6)
6.1. Uncertainty Quantification Performance
RQ1 evaluates the performance of $\text{R}^{2}$C compared to other UQ methods. Table 1 presents results on five RAR systems across three datasets. The findings indicate that most white-box methods, i.e., PE (Kadavath and others, 2022), SE (Kuhn et al., 2023), MARS (Bakman et al., 2024), and SAR (Duan et al., 2024), perform relatively poorly, with AUROC values ranging from about 30 to 60. This weakness stems from their overreliance on token probabilities. In contrast, black-box methods, such as NumSS (Kuhn et al., 2023), EigV, ECC, and Deg (Lin et al., 2024), generally outperform white-box methods, reaching average AUROC values between roughly 60 and 80. Interestingly, P(true) ranks as the second-best method in terms of average AUROC across all approaches, highlighting that black-box methods become much more effective when using an LLM as the scoring function. This advantage largely comes from their stronger reliance on textual diversity. The supervised method LARS, optimized for QA with no reasoning, achieves AUROC scores between 65 and 85 across all cases, surpassing both white-box and black-box approaches on average. This finding highlights the potential of as a promising direction for future research.
@@ 6.2. Extrinsic Evaluation Results
For model selection, Table 3 reports the results of individual RAG and RAR systems, as well as selection-based ensemble models. Among individual systems, RAR models outperform RAG models for both simpler single-hop questions (PopQA) and more complex multi-hop queries (HotpotQA and Musique). Within selection-based ensemble models, LLMBLender performs poorly, even worse than fine-tuned single RAR models such as ReSearch and Search-R1. In contrast, RAGEnsemble surpasses both LLMBLender and all single RAR models, highlighting the advantage of an instruction-tuned model for selecting the final response. Finally, $\text{R}^{2}$C Select achieves the best performance, significantly outperforming all RAR and RAG systems, with 3.7% average improvement on HotpotQA and Musique datasets. These results confirm that $\text{R}^{2}$C score is a reliable criterion, not only within a single system but also across different systems. Interestingly, using $\text{R}^{2}$C confidence scores alone, without clustering, improves selection-based ensemble models in comparable settings, highlighting informativeness of $\text{R}^{2}$C scores even for models that are not trained on confidence.
Figure 6. Performance of $\text{R}^{2}$C with different action sets.
Table 4. Performance of UQ methods measured by AUROC across different LLMs. For each LLM and for each column, the best and second-best methods are indicated by bold and underline, respectively. Superscripts and denote statistically significant differences according to the DeLong test ($p<0.05$), compared to SelfC and ReaC, respectively, which are the two best-performing methods on average.
Gemma2-9b-it GPT-4o-mini
RAG SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) Avg. SelfAsk (Press et al., 2023) ReAct (Yao et al., 2023) Search-o1 (Li et al., 2025b) Avg.
Uncer. M. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq. Popqa Hotpot Musiq.
RrrC (Li et al., 2025a) 76.64 71.01 76.16 57.74 72.93 75.32 55.72 63.96 64.33 68.20 63.07 70.96 83.69 53.21 58.16 54.78 70.49 77.27 84.04 68.41
P(true) (Kadavath and others, 2022) 71.33 74.90 73.44 76.56 73.08 75.26 76.32 72.45 74.62 74.22
SelfC (Wang et al., 2023) 76.82 72.70 73.59 81.79 74.34 76.84 79.17 70.59 76.07 75.77 76.59 58.28 54.04 81.08 73.32 75.83 78.52 74.39 81.83 72.65
ReaC (Qi et al., 2025) 76.24 79.77 78.73 80.16 71.33 80.80 78.45 73.61 75.77 77.21 75.53 60.58 54.66 81.53 74.34 81.10 79.41 75.09 85.09 74.15
$\text{R}^{2}$C (our) 80.29 81.86 80.31 82.87 80.62 80.37 87.65 76.95 77.39 80.92 86.56 87.82 84.28 81.56 75.84 81.94 82.48 84.59 86.29 83.48
@@ 6.3. Explaining the Strength of $\text{R}^{2}$C
RQ3 examines why $\text{R}^{2}$C outperforms other UQ methods. To this end, we analyze the diversity of reasoning paths as an indicator of how effectively $\text{R}^{2}$C captures the uncertainty arising from both the generator and the retriever, i.e., the two key sources of uncertainty in our framework. We report the number of unique documents retrieved during the multi-generation step, as shown in Figure 2. On average, $\text{R}^{2}$C retrieves 24.71 unique documents for a single confidence score, whereas RrrC, SelfC, and ReaC retrieve 5.81, 15.35, and 16.4 documents, respectively.
@@
where $cos(.)$ represents cosine similarity function, and $e_{i}$ denotes the normalized embedding of query $q_{i}$ (i.e., $|e_{i}|=1$) obtained from sentence-transformers/all-MiniLM-L6-v2. Figure 3 presents the results. On average, $\text{R}^{2}$C achieves a diversity score of $0.35$, while RrrC, SelfC, and ReaC achieve $0.20$, $0.28$, and $0.30$, respectively. These results indicate that $\text{R}^{2}$C generates more diverse search queries when estimating confidence. These findings indicate $\text{R}^{2}$C effectively captures the uncertainty of both the retriever and the generator by sequentially diversifying their inputs.
@@
For a deeper exploration of the trade-offs between effectiveness and efficiency, we measure efficiency using the total number of generated tokens, following prior work (Perez-Beltrachini and Lapata, 2025). Figure 5 reports AUROC performance for equal numbers of generated tokens. It shows $\text{R}^{2}$C achieves higher AUROC scores than other methods given the same number of generated tokens for multi-hop datasets. For the single-hop dataset, PopQA, $\text{R}^{2}$C performs comparably to SelfAsk, ReAct, and ReSearch, but surpasses them on Search-o1 and Search-R1. Moreover, $\text{R}^{2}$C produces on average around 700 tokens with three generations, reaching the same AUROC score of baselines with approximately 1,700 tokens with 10 generations. These findings demonstrate that $\text{R}^{2}$C improves efficiency by roughly 2.5 times. Overall, these results indicate that $\text{R}^{2}$C outperforms other UQ methods in terms of both effectiveness and token generation efficiency. Beyond improving token-generation efficiency, $\text{R}^{2}$C is well suited for real-world deployment because it enables parallel generation of fully independent responses, thereby addressing the typical latency concerns associated with sampling-based methods. In practice, the confidence score can be computed at the cost of at most one additional generation beyond the most-likely response.